# Investigating the Upper Scorpius OB association with HERMES. I. The spectroscopic sample and 6D kinematics

### v1

@@ Abstract
OB associations are large, unbound groups of young stars, which typically exhibit great complexity in spatial, kinematic and age structure, which hints at formation scenarios involving an intricate interplay of molecular cloud turbulent kinematics and stellar feedback over extended periods. The kinematic properties of the numerous low-mass populations within OB associations can provide valuable constraints on their initial configurations, and thus the dominant mechanisms driving star formation and their dispersal into the field. However, precise kinematic information and accurate membership determination is necessary to build a complete and detailed description of OB associations, to test the many competing hypotheses for their formation. We present results from a large spectroscopic survey of the nearest young association to the Sun, Upper Scorpius, conducted using 2dF/HERMES on the Anglo-Australian Telescope. We use spectroscopic youth criteria such as Li-equivalent widths to identify 1000 pre-main sequence (PMS) members across the region and measure radial velocities, combining these with Gaia EDR3 5-parameter astrometry to obtain 6D kinematic information. We separate confirmed PMS association members into distinct kinematic groups and measure expansion and rotation trends in each. We find strong evidence for asymmetric expansion in several groups and derive expansion timescales from the greatest rates of expansion in each group. We also trace the past motion of these groups using an epicycle approximation and estimate the time since their most compact configuration. These kinematic properties are compared to literature ages and the star formation history of Upper Scorpius is discussed. We find evidence that a scenario in which star formation in the subgroups of Upper Scorpius proceeded independently, either by self-instability or external feedback from Upper Centaurus-Lupus, is more likely than a recently proposed “cluster chain” scenario in which these subgroups have triggered each other.
@@
However, it has yet to be determined what are the dominant mechanisms driving star formation in these complexes. It has been suggested that the age gradients found in OB associations point toward multiple epochs of star formation, with the initiation of star formation in one epoch being triggered by the impact of feedback from massive stars formed in the previous epoch (e.g., Elmegreen & Lada, 1977; Posch et al., 2025; Kerr et al., 2025). There is also the suggestion that star formation could be initiated by collisions between giant molecular clouds (GMCs) (e.g., Scoville et al., 1986), with a model in which collisions are driven by galactic shearing motions of orbiting GMCs (Tan, 2000; Tasker & Tan, 2009; Li et al., 2018) being potentially relevant to explaining observed large-scale star formation properties of disk galaxies (Tan, 2010; Suwannajak et al., 2014; Fortune-Bashee et al., 2024). Simulations of star formation resulting from individual GMC-GMC collisions, with relative cloud velocities of can produce spatial structure and disturbed kinematics among the newly formed stars similar to those observed in nearby regions (e.g., Wu et al., 2017, 2020). Stellar feedback driven cloud collisions have also been proposed as the main driver of new star formation in galaxies (Inutsuka et al., 2015). Other possibilities include models of star formation regulated by turbulence in GMCs (e.g., Krumholz & McKee, 2005; Padoan & Nordlund, 2011; Hennebelle & Chabrier, 2011). Combinations of the above models are also possible, e.g., if shear-driven GMC collisions are the main driver of turbulence in GMCs (Tan et al., 2013; Wu et al., 2018).
@@ 2.1 AAT observations
Observations were made on April 16th, May 21st - 22nd & June 10th - 12th 2023 with the 2-degree field (Lewis et al., 2002, 2dF) fibre positioner and the high-efficiency and resolution multi-element spectrograph (Sheinis et al., 2015, HERMES), which provides a resolution of 28,000 in four optical bands, at the Anglo-Australian Telescope (AAT). The red band of HERMES covers the wavelength range 6478–6737 Å within which both the H 6562.8 Å and Li 6707.8 Å lines are found. For each field we selected targets based on colour-absolute magnitude diagram (CaMD) position, which totaled 8061 targets for all fields with overlap, 7039 of which were unique sources. Multiple 1800s exposure were taken for each field, the total time for which, along with numbers of targets and numbers of spectroscopically confirmed PMS stars (Section 2.1.3), are given in Table 1. Calibration frames, including dark frames and multifibre flat fields were taken for each field and 25-40 fibres were used to measure the sky spectrum for each field.
@@ 2.1.2 Radial Velocities
Radial velocities are the last component of complete 6D position-velocity information missing from Gaia for most sources. With high-resolution spectra from HERMES we can measure radial velocities up to a precision of , which is sufficient to investigate the internal kinematics of young clusters and star-forming regions.
@@ 2.2 Gaia astrometry
Spectroscopic targets are matched to the Gaia DR3 catalogue (Gaia Collaboration et al., 2021) to obtain positions, proper motions and parallaxes. Combined with our RVs, this gives us full 6D kinematic information for all of our 1204 confirmed YSOs. We filter targets with poor Gaia astrometric quality following the recommended criteria of Fabricius et al. (2021); RUWE 1.4, ipd_frac_multi_peak 3 and ipd_gof_harmonic_amplitude 0.1, in order to remove non-single stars. The velocity contribution of binary orbits, in particular, could bias our kinematic analysis. After applying these astrometric quality filters we are left with 544 confirmed YSOs with quality 6D kinematic information.
@@ 2.3 Deriving 3D positions and velocities
When deriving 3D positions we use Gaia parallaxes rather than distance estimates from Bailer-Jones et al. (2021). As described in Fiorellino et al. (2024), the priors involved in the probabilistic estimation of distances by Bailer-Jones et al. (2021) are not optimised for YSOs, nor for sources in clustered regions.
@@ 2.4 Summary of the data
After reducing the spectra of our AAT sources, we have 6646 unique sources with spectroscopic RV and EW(Li) values spread across 25 fields over the Upper Sco region, with 5-parameter astrometry from Gaia EDR3, with which we calculate Cartesian XYZ positions and UVW velocities. 1204 of these sources are identified as YSOs with significant EW(Li)s, EW(H)s or Gaia variability flag. The median uncertainties on parallax, distance, proper motion and RV for the confirmed PMS stars are 0.05 mas, 0.9 pc, 0.05 mas yr-1 and 0.56 km s-1 respectively. 544 of these confirmed PMS stars pass our filters for astrometric quality.

### v2

@@ Abstract
OB associations are large, unbound groups of young stars, which typically exhibit great complexity in spatial, kinematic and age structure, hinting at formation scenarios involving an intricate interplay of molecular cloud turbulent kinematics and stellar feedback over extended periods. The kinematic properties of the numerous low-mass populations within OB associations can provide valuable constraints on their initial configurations, and thus the dominant mechanisms driving star formation and their dispersal into the field. We present results from a large spectroscopic survey of the nearest young association to the Sun, Upper Scorpius, conducted using 2dF/HERMES on the Anglo-Australian Telescope. We use spectroscopic youth criteria such as Li-equivalent widths to identify 1000 pre-main sequence (PMS) members across the region and measure radial velocities, combining these with Gaia EDR3 5-parameter astrometry to obtain 6D kinematic information. We separate confirmed PMS association members into distinct kinematic groups and measure expansion and rotation trends in each. We also trace the past motion of these groups using an epicycle approximation and estimate the time since their most compact configuration. These kinematic properties are compared to literature ages and the star formation history of Upper Scorpius is discussed. We find evidence that a scenario in which star formation in the subgroups of Upper Scorpius proceeded independently, either by self-instability or external feedback from Upper Centaurus-Lupus, is more likely than a recently proposed “cluster chain” scenario in which these subgroups have triggered each other.
@@
However, it has yet to be determined what are the dominant mechanisms driving star formation in these complexes. It has been suggested that the age gradients found in OB associations point toward multiple epochs of star formation, with the initiation of star formation in one epoch being triggered by the impact of feedback from massive stars formed in the previous epoch (e.g., Elmegreen & Lada, 1977; Posch et al., 2025; Kerr et al., 2025). There is also the suggestion that star formation could be initiated by collisions between giant molecular clouds (GMCs) (e.g., Scoville et al., 1986), with a model in which collisions are driven by galactic shearing motions of orbiting GMCs (Tan, 2000; Tasker & Tan, 2009; Li et al., 2018) being potentially relevant to explaining observed large-scale star formation properties of disk galaxies (Tan, 2010; Suwannajak et al., 2014; Fortune-Bashee et al., 2024). Simulations of star formation resulting from individual GMC-GMC collisions, with relative cloud velocities of can produce spatial structure and disturbed kinematics among the newly formed stars similar to those observed in nearby regions (e.g., Wu et al., 2017; Wu et al., 2020). Stellar feedback driven cloud collisions have also been proposed as the main driver of new star formation in galaxies (Inutsuka et al., 2015). Other possibilities include models of star formation regulated by turbulence in GMCs (e.g., Krumholz & McKee, 2005; Padoan & Nordlund, 2011; Hennebelle & Chabrier, 2011). Combinations of the above models are also possible, e.g., if shear-driven GMC collisions are the main driver of turbulence in GMCs (Tan et al., 2013; Wu et al., 2018).
@@ 2.1 AAT observations
Observations were made on April 16th, May 21st - 22nd & June 10th - 12th 2023 with the 2-degree field (Lewis et al., 2002, 2dF) fibre positioner and the high-efficiency and resolution multi-element spectrograph (Sheinis et al., 2015, HERMES), which provides a resolution of 28,000 in four optical bands, at the Anglo-Australian Telescope (AAT) as part of program 2023A/30. The red band of HERMES covers the wavelength range 6478–6737 Å within which both the H 6562.8 Å and Li 6707.8 Å lines are found. For each field we selected targets based on colour-absolute magnitude diagram (CaMD) position, which totaled 8061 targets for all fields with overlap, 7039 of which were unique sources. Multiple 1800s exposure were taken for each field, the total time for which, along with numbers of targets and numbers of spectroscopically confirmed PMS stars (Section 2.1.3), are given in Table 1. Calibration frames, including dark frames and multifibre flat fields were taken for each field and 25-40 fibres were used to measure the sky spectrum for each field.
@@ 2.1.2 Radial Velocities
Radial velocities are the last component of complete 6D position-velocity information missing from Gaia for most sources. With high-resolution spectra from HERMES we can measure radial velocities up to a typical precision of (Zwitter et al., 2018), which is sufficient to investigate the internal kinematics of young clusters and star-forming regions.
@@ 2.2 Gaia astrometry
Spectroscopic targets are matched to the Gaia DR3 catalogue (Gaia Collaboration et al., 2021) to obtain positions, proper motions and parallaxes. Combined with our RVs, this gives us full 6D kinematic information for all of our 1204 confirmed YSOs. We filter targets with poor Gaia astrometric quality following the recommended criteria of Fabricius et al. (2021); RUWE 1.4, ipd_frac_multi_peak 3 and ipd_gof_harmonic_amplitude 0.1, in order to remove non-single stars. The velocity contribution of binary orbits, in particular, could bias our kinematic analysis. We also remove targets with RV km s-1 or with RV uncertainty km s-1, in order to further filter sources with unreliable RVs and possible binaries. Out of our 1204 confirmed YSOs, 79 fail the ipd_gof_harmonic_amplitude criterion, 92 fail the ipd_frac_multi_peak criterion, 251 fail the RUWE criterion and 482 fail the RV criteria, with overlap. After applying these astrometric quality filters we are left with 544 confirmed YSOs with quality 6D kinematic information.
@@ 2.3 Deriving 3D positions and velocities
When deriving 3D positions we use Gaia parallaxes rather than distance estimates from Bailer-Jones et al. (2021). As described in Fiorellino et al. (2024), the priors involved in the probabilistic estimation of distances by Bailer-Jones et al. (2021) are not optimised for YSOs, nor for sources in clustered regions. We apply the parallax zero point correction to our sources, as prescribed in Lindegren et al. (2021), as a function of ecliptic latitude, magnitude and colour. We also apply the magnitude dependent correction to parallax uncertainties prescribed by El-Badry et al. (2021).
@@ 2.4 Summary of the data
After reducing the spectra of our AAT sources, we have 6646 unique sources with spectroscopic RV and EW(Li) values spread across 25 fields over the Upper Sco region, with 5-parameter astrometry from Gaia EDR3, with which we calculate Cartesian XYZ positions and UVW velocities. 1204 of these sources are identified as YSOs with significant EW(Li)s, EW(H)s or Gaia variability flag. The median uncertainties on parallax, distance, proper motion and RV for the confirmed PMS stars are 0.05 mas, 0.9 pc, 0.05 mas yr-1 and 0.56 km s-1 respectively, with a median SNR of their HERMES spectra. 544 of these confirmed PMS stars pass our filters for astrometric quality.